OpenStudy (anonymous):

One fourth of a number is two more than one fifth of the number. What is the number?

OpenStudy (anonymous):

you convert this to an equation: "x" being your number 1.4x = 2 + 1/5x and solve accordingly~

OpenStudy (anonymous):

(1/4)x=2+(1/5)x (1/4)x-(1/5)x=2 (5/20)x-(4/20)x=2 (1/20)x=2 x=40 Check: (1/4)40=2+(1/5)40 10=2+8 10=10

OpenStudy (anonymous):

1/4x=2+1/5x

OpenStudy (anonymous):

To make the math easier, give 1/4x and 1/5x the same divisor: 5/20x = 2 + 4/20x Then just solve 1/20x = 2 x = 1/20 (2) = 2/20 = 10 so your number "x" = 10